OpenStudy (anonymous):

Calculate the density of a substance in pen- nyweights per peck that has a volume of 0.333 hogsheads and a mass of 49.62 kg. 1 pennyweight = 1.55 grams 1 hogshead = 238.5 liters 1 peck = 9.091 liters. Answer in units of pw/pk

OpenStudy (anonymous):

please helppp

OpenStudy (paxpolaris):

\[{49.62\ \cancel {kg} \over0.333\ \cancel {hogsheads}}\times {1000\ g \over \cancel {kg}} \times {1\ pennyweight \over 1.55\ g} \times {1\ \cancel {hogshead}\over 238.5\ l} \times {9.091\ l \over 1\ peck}\]

OpenStudy (anonymous):

is it 5007768759.091 ?